I need help to find the intersection point of the graphs of the given equations:
y=10x-8
y=3x+2

Answer by stanbon(75887)

I need help to find the intersection point of the graphs of the given equations:
y=10x-8
y=3x+2
-----
Equation:
10x-8 = 3x+2
7x = 10
x = 10/7
=============
Solve for "y";
y = 3x+2
y = 3(10/7) + (14/7)
y = 44/7
====================
Cheers,
Stan H.
====================